# Metric Pairs and Tuples in Theory and Applications

### v1

@@ 5.2. Simplicial Hausdorff Distance
In [nnadi], Nnadi and Isaksen defined a version of the Hausdorff distance between simplicial complexes and, moreover, between simplicial filtrations. To do that, they consider the class $\mathbb{X}^{d}$ of pairs $(X,f)$, where $X$ is a finite simplicial complex, $f\colon X_{0}\to\mathbb{R}^{d}$ is an injective map and $X_{0}$ is the set of vertices of $X$. They also consider simplicial filtrations as sets $\{X_{\alpha}\}$ of simplicial complexes such that $X_{\alpha}\subseteq X_{\beta}$ for $\alpha\leq\beta$.
@@
where $\operatorname{Conv}(\cdot)$ denotes the convex hull. It is not difficult to see that
@@
where $\delta$ denotes the simplicial Hausdorff distance from [nnadi].
@@
whereas, by definition,
@@
Furthermore, we observe that it is possible to globalize the distance between simplicial complexes introduced in [nnadi] à la Gromov, as follows: consider pairs of the form $(X,\mathcal{X})$ where $X$ is a finite metric space and $\mathcal{X}$ is a simplicial complex with set of vertices $X$. Then, given two such pairs, $(X,\mathcal{X})$ and $(Y,\mathcal{Y})$, we can define their simplicial Gromov–Hausdorff distance by
@@
where the infimum runs over metric spaces $Z$ and isometric embeddings $\phi\colon X\to Z$, $\psi\colon Y\to Z$, and where $\delta^{Z}$ is analogous to the simplicial Hausdorff distance, but for simplicial complexes embedded in $Z$.
It is obvious that $\delta_{\operatorname{GH}}$ is non-negative and symmetric. Moreover, the triangle inequality follows from the triangle inequality satisfied by the simplicial Hausdorff distance (which can be proved in any metric space $Z$ just as in $\mathbb{R}^{d}$) and an argument analogous to the proof of [jansen2017, Proposition 1.6].
Finally, we say that $(X,\mathcal{X})$ and $(Y,\mathcal{Y})$ are isomorphic if there exists an isometry $\phi\colon X\to Y$ such that $\sigma$ is a $k$-simplex in $\mathcal{X}$ if and only if $\phi(\sigma)$ is a $k$-simplicial complex in $\mathcal{Y}$. In that case, by taking $Z=Y$ and $\psi=\mathrm{id}_{Y}$, it easily follows that $\delta^{Z}((\mathcal{X},\phi),(\mathcal{Y},\psi))=0$, which in turn implies $\delta_{\operatorname{GH}}((X,\mathcal{X}),(Y,\mathcal{Y}))=0$. Conversely, if $\delta_{\operatorname{GH}}((X,\mathcal{X}),(Y,\mathcal{Y}))=0$ then (following again the proof of [jansen2017, Proposition 1.6]), for any $n\in\mathbb{N}$ there exist metric spaces $Z_{n}$ and isometric embeddings $\phi_{n}\colon X\to Z_{n}$, $\psi_{n}\colon Y\to Z_{n}$ such that
@@
We could even extend $\delta_{\operatorname{GH}}$ to simplicial filtrations in a natural way and by (5.2) and (5.3), an inequality of the form
$\hat{\delta}_{\operatorname{GH}}((X,\mathcal{X}),(Y,\mathcal{Y}))\leq d_{\operatorname{GH}}(\{X_{\alpha}\},\{Y_{\alpha}\})\leq(k+1)\hat{\delta}_{\operatorname{GH}}((X,\mathcal{X}),(Y,\mathcal{Y}))$
follows. We leave the details to the reader.
@@ 5.3. Persistence Matching Diagrams
Persistent homology is a fundamental tool in Topological Data Analysis (TDA). Its application to filtrations of simplicial complexes yields barcodes or persistence diagrams, where intervals or points represent the birth and death of homological features across the filtration. In the following sections, we assume the reader is familiar with basic notions of TDA. For those less acquainted with the subject, we recommend the introductory references [munch2017user, virk2022introduction].
@@
González-Díaz, Soriano-Trigueros and Torras-Casas presented in [torras_rocio_2] a particular block function $\mathcal{M}_{f}$ to obtain partial matchings between persistence diagrams induced by a morphism $f$. In [torrascasas2024stability], González-Díaz and Torras-Casas gave stability results for the persistence matching diagrams based on a block function $\mathcal{M}_{f}^{0}$, where the morphism $f$ is the inclusion of $X\subseteq Z$ and they pay attention to the 00–homology barcode.
@@
$\tilde{d}_{\operatorname{GH}}((Z,X),(Z’,X’))\leq d_{\operatorname{GH}}((Z,X),(Z’,X’))\leq 2\ \tilde{d}_{\operatorname{GH}}((Z,X),(Z’,X’))$ ’ ) )

### v2

@@ 5.2. Simplicial Hausdorff Distance
In [undefan], Nnadi and Isaksen defined a version of the Hausdorff distance between simplicial complexes. To do that, they consider the class $\mathbb{X}^{d}$ of pairs $(X,f)$, where $X$ is a finite simplicial complex, $f\colon X_{0}\to\mathbb{R}^{d}$ is an injective map and $X_{0}$ is the set of vertices of $X$, and defined the distance as follows:
Definition 5.1 (Definition 2.2, [undefan]).
The simplicial Hausdorff distance is a map $\delta\colon\mathbb{X}^{d}\times\mathbb{X}^{d}\to\mathbb{R}_{\geq 0}$, such that
(5.1) $\delta\left((X,f),(Y,g)\right)=\max\left\{\vec{d}((X,f),(Y,g)),\vec{d}((Y,g),(X,f))\right\},$
where the directed distance $\vec{d}((X,f),(Y,g))=\inf\{\epsilon>0\colon(X,f)\text{ is }\epsilon\text{-close to }(Y,g)\}$. The pair $(X,f)$ is $\epsilon$-close to $(Y,g)$ if and only if: for every $k$-simplex $\sigma$ in $X$, there is a $k$-simplex $\tau$ in $Y$ such that for any vertex $v\in\sigma$, there is some vertex $w\in\tau$ with $d(f(v),g(w))<\epsilon$, for every $k$.
Moreover, the authors also defined a distance between simplicial filtrations. Consider simplicial filtrations as sets $\{X_{\alpha}\}$ of simplicial complexes such that $X_{\alpha}\subseteq X_{\beta}$ for $\alpha\leq\beta$. Additionally, they also denote $(X,f)$ as a filtered complex, such that for every $\alpha>0$, $X_{\alpha}$ is a simplicial subcomplex on the vertex set $X_{0}$ with index value $\alpha$ and $f\colon X_{0}\to\mathbb{R}^{d}$ a measurement function on the vertex set. Then
Definition 5.2 (Definition 2.20, [undefan]).
The simplicial Hausdorff distance between two filtered complexes $(X,f)$ and $(Y,g)$ is
(5.2) $\hat{\delta}\left((X,f),(Y,g)\right)=\max\left\{\vec{\mathrm{d}}((X,f),(Y,g)),\vec{\mathrm{d}}((Y,g),(X,f))\right\},$
where $\vec{\mathrm{d}}((X,f),(Y,g))=\sup_{\alpha>0}\{\vec{d}((X_{\alpha},f),(Y_{\alpha},g))\}$.
@@
where $\operatorname{Conv}(\cdot)$ denotes the convex hull, and $D\in\mathbb{N}$ big enough such that $|(X,f)|$ can be correctly embedded, i.e., $D\geq|X_{0}|$ would be enough. It is not difficult to see that
@@
where $\delta$ denotes the simplicial Hausdorff distance from [undefan].
@@
where $\hat{\delta}$ denotes the simplicial Hausdorff distance between filtered complexes, as defined in Definition 5.2. On the other hand, by definition,
@@
Furthermore, we observe that it is possible to globalize the distance between simplicial complexes introduced in [undefan] à la Gromov, as follows: consider pairs of the form $(X,\mathcal{X})$ where $X$ is a finite metric space and $\mathcal{X}$ is a simplicial complex with set of vertices $X$. Then, given two such pairs, $(X,\mathcal{X})$ and $(Y,\mathcal{Y})$, we can define their simplicial Gromov–Hausdorff distance by
@@
where the infimum runs over metric spaces $(Z,d_{Z})$ and isometric embeddings $\phi\colon X\to Z$, $\psi\colon Y\to Z$, and where $\delta^{Z}$ is analogous to the simplicial Hausdorff distance, but for simplicial complexes embedded in $Z$, i.e.
$\delta^{Z}((\mathcal{X},\phi),(\mathcal{Y},\psi))=\max\left\{\vec{d}_{Z}((\mathcal{X},\phi),(\mathcal{Y},\psi)),\vec{d}_{Z}((\mathcal{Y},\psi),(\mathcal{X},\phi))\right\}$
where
$\vec{d}_{Z}((\mathcal{X},\phi),(\mathcal{Y},\psi))=\max_{k}\max_{\begin{subarray}{c}\sigma\in\mathcal{X}\\ \dim(\sigma)=k\end{subarray}}\min_{\begin{subarray}{c}\tau\in\mathcal{Y}\\ \dim(\tau)=k\end{subarray}}\max_{v\in\sigma}\min_{w\in\tau}d_{Z}(\phi(v),\psi(w)).$
It is obvious that $\delta_{\operatorname{GH}}$ is non-negative and symmetric. Moreover, the triangle inequality follows from the triangle inequality satisfied by the simplicial Hausdorff distance (which can be proved in any metric space $Z$ just as in $\mathbb{R}^{d}$) and an argument analogous to the proof of [undefah, Proposition 1.6].
Finally, we say that $(X,\mathcal{X})$ and $(Y,\mathcal{Y})$ are isomorphic if there exists an isometry $\phi\colon X\to Y$ such that $\sigma$ is a $k$-simplex in $\mathcal{X}$ if and only if $\phi(\sigma)$ is a $k$-simplicial complex in $\mathcal{Y}$. In that case, by taking $Z=Y$ and $\psi=\mathrm{id}_{Y}$, it easily follows that $\delta^{Z}((\mathcal{X},\phi),(\mathcal{Y},\psi))=0$, which in turn implies $\delta_{\operatorname{GH}}((X,\mathcal{X}),(Y,\mathcal{Y}))=0$. Conversely, if $\delta_{\operatorname{GH}}((X,\mathcal{X}),(Y,\mathcal{Y}))=0$ then (following again the proof of [undefah, Proposition 1.6]), for any $n\in\mathbb{N}$ there exist metric spaces $Z_{n}$ and isometric embeddings $\phi_{n}\colon X\to Z_{n}$, $\psi_{n}\colon Y\to Z_{n}$ such that
@@
We could even extend $\delta_{\operatorname{GH}}$ to simplicial filtrations: given simplicial filtrations $(X,\mathcal{X})=\{(X_{\alpha},\mathcal{X}_{\alpha})\}$ and $(Y,\mathcal{Y})=\{(Y_{\alpha},\mathcal{Y}_{\alpha})\}$, let
$\hat{\delta}_{\operatorname{GH}}((X,\mathcal{X}),(Y,\mathcal{Y}))=\max\left\{\vec{d}_{Z}((\mathcal{X},\phi),(\mathcal{Y},\psi)),\vec{d}_{Z}((\mathcal{Y},\psi),(\mathcal{X},\phi))\right\}$
where
$\vec{d}_{Z}((\mathcal{X},\phi),(\mathcal{Y},\psi))=\sup_{\alpha}\vec{d}_{Z}((\mathcal{X}_{\alpha},\phi),(\mathcal{Y}_{\alpha},\psi)).$
By (5.4) and (5.5), the equivalence inequality follows:
$\hat{\delta}_{\operatorname{GH}}((X,\mathcal{X}),(Y,\mathcal{Y}))\leq d_{\operatorname{GH}}(\{X_{\alpha}\},\{Y_{\alpha}\})\leq(k+1)\hat{\delta}_{\operatorname{GH}}((X,\mathcal{X}),(Y,\mathcal{Y})).$
@@ 5.3. Persistence Matching Diagrams
Persistent homology is a fundamental tool in Topological Data Analysis (TDA). Its application to filtrations of simplicial complexes yields barcodes or persistence diagrams, where intervals or points represent the birth and death of homological features across the filtration. In the following sections, we assume the reader is familiar with basic notions of TDA. For those less acquainted with the subject, we recommend the introductory references [undefal, undefav].
@@
González-Díaz, Soriano-Trigueros and Torras-Casas presented in [undefz] a particular block function $\mathcal{M}_{f}$ to obtain partial matchings between persistence diagrams induced by a morphism $f$. In [undefau], González-Díaz and Torras-Casas gave stability results for the persistence matching diagrams based on a block function $\mathcal{M}_{f}^{0}$, where the morphism $f$ is the inclusion of $X\subseteq Z$ and they pay attention to the 0–homology barcode.
@@
$\tilde{d}_{\operatorname{GH}}((Z,X),(Z^{\prime},X^{\prime}))\leq d_{\operatorname{GH}}((Z,X),(Z^{\prime},X^{\prime}))\leq 2\ \tilde{d}_{\operatorname{GH}}((Z,X),(Z^{\prime},X^{\prime}))$